OpenStudy (anonymous):

Chris eats 3 slices of a pizza that is cut into 8 equal slices. What percentage of the pizza does he eat?

Parth (parthkohli):

=> 3/8 as a fraction. => Multiply by 100 and you get \[\LARGE {3.100 \over 8} = {3.25 \over 2} = {75\over 2} = 37.5percent\]

OpenStudy (anonymous):

I think I would approach it as a ratio. \[3/8=x/100\] Find x for either using algebra or cross multiplying and then dividing. Using algebra, use get the x on one side of the equation by multiplying both sides by 100. So that gives you 100(3/8)=x Use your calculator to divide 8 into 3 to get .375 and then multiply by 100 (.375)(100)= 37.5 percent